OpenStudy (anonymous):

What is the result of adding the system of equations? 2x + y = 4 3x - y = 6 a. x = 2 b. x = 10 c. 5x = 10

OpenStudy (anonymous):

i think its c

OpenStudy (hyper):

Yea, its c

OpenStudy (hyper):

Solution: (2x +y)+(3x -y) = 4 + 6 5x = 10

OpenStudy (anonymous):

thank you so much!! =))